OpenStudy (misssmartiez):

What is the mean of absolute deviation (MAD) of the following data set 2, 2, 5 A) 0 B) 2/3 C) 4/3 D) 1/3

Vocaloid (vocaloid):

First step: what is the mean?

OpenStudy (misssmartiez):

9.

OpenStudy (misssmartiez):

@sweetburger @rebeccaxhawaii

Vocaloid (vocaloid):

nope, the mean is not 9. add the numbers together, then divide by 3

OpenStudy (misssmartiez):

OHHH THAT IS WHAT I LACKED IN MY STEPS!

Vocaloid (vocaloid):

cool, let me know what you get as your final answer

OpenStudy (misssmartiez):

@Vocaloid

OpenStudy (misssmartiez):

3 - 2 = 1. 3 - 2= 1 3- 5 = l-2l = 2 1 + 1 + 2 = 4 / 3 = 1.3333(continue) 1.33333 as fraction = 4/3

Vocaloid (vocaloid):

good!